# Topological Weyl Altermagnetism in CrSb

### v1

@@ Abstract
Altermagnets constitute a novel, third fundamental class of collinear magnetic ordered materials, alongside with ferro- and antiferromagnets. They share with conventional antiferromagnets the feature of a vanishing net magnetization. At the same time they show a spin-splitting of electronic bands, just as in ferromagnets, caused by the atomic exchange interaction. Interestingly, the altermagnetic spin-splitting may in principle allow for topological electronic structures far beyond collinear antiferromagnets with spin-degenerate bands. Here, using high-resolution and spin angle-resolved photoemission spectroscopy, we observe that the momentum-dependent spin-splitting in altermagnetic CrSb reaches up to 1 eV, while vanishing at high symmetry lines in the Brillouin zone. We find it causes spin-carrying altermagnetic Weyl nodes, and observe their protected Fermi-arc surface states. This establishes that in altermagnets the large energy scale intrinsic to the spin-splitting – orders of magnitude larger that the relativistic spin-orbit coupling – creates its own realm of robust electronic topology, with spin-polarized topological surface states that are altermagnetic.
@@
The recently established new class of altermagnetic materials is characterized by the presence of particular alternating opposite-spin sublattices. They defy conventional magnetic classifications by breaking time-reversal symmetry while simultaneously maintaining a zero net magnetization enforced by spin-lattice symmetry[1, 2, 3]. Precisely this intrinsic trait allows altermagnets to combine advantageous features of both ferromagnets and antiferromagnets[4, 5], including spin-transport properties and compatibility with a diverse range of materials[6, 7, 8, 9, 10, 11, 12, 13], from superconductors[14, 15, 17, 18, 16] to insulators[1, 2, 19, 20, 21]. As the vanishing net magnetization in altermagnets avoids effects of long-range magnetic stray fields typical for ferromagnets, this combination opens up promising avenues for applications in magneto-optics[22], spintronics[24, 23, 9], and beyond.
@@
Interestingly, symmetry analysis reveals that the electronic structure of altermagnets may in principle provide fertile ground for non-trivial electronic topology, due to the unique interplay between time-reversal and crystalline symmetries[26, 25, 27, 28]. Despite extensive experimental and theoretical scrutiny of altermagnets[1, 2, 20, 29, 19, 21, 30, 31, 32] and topological semimetals[33, 34, 35, 36, 37, 38, 39, 40, 34, 35] separately, the simultaneous experimental manifestation of their characteristic phenomena has remained elusive so far. Here we establish CrSb as a topological Weyl semimetal with two distinct types of spin-carrying nodal structures. The giant spin splitting that we find from high resolution angle-resolved photoemission spectroscopy (ARPES) and spin-ARPES measurements aligns well with density functional theory (DFT) calculations on CrSb, as do the bulk Weyl points (WPs) with linear dispersion and ensuing altermagnetic surface Fermi arcs (SFAs) that connect same-spin WPs at the surface.
@@
For both its altermagnetic and topological features, the symmetries of CrSb are crucial. Figure 1a shows the crystal and magnetic structure of CrSb, which forms a hexagonal structure with the space group $P6_{3}/mmc$ (No. 194)[41]. The magnetic space group is $P6^{\prime}_{3}/m^{\prime}m^{\prime}c$ (No. 194.268) and the spin space group is $P^{-1}{6_{3}/}^{-1}{m}^{1}{m}^{-1}c^{\infty m}1$ with a six-fold screw rotation symmetry connecting Cr-atoms of opposite magnetic moment in real space. The three-dimensional (3D) Fermi surface of CrSb from DFT calculations in Fig. 1c illustrates that sixfold rotation connecting opposite-spin sublattices in real space (Fig. 1a) also connects opposite-spin electronic states in momentum space (Fig. 1c), as is typical for altermagnets[1, 2]. The calculated Fermi surface also illustrates the four mirror planes in the Brillouin zone (BZ) on which the altermagnetic bands are spin-degenerate in absence of spin-orbit coupling (SOC), as shown in Fig. 1b.
@@
The altermagnetic spin-splitting allows for topological electronic structures far beyond antiferromagnets with completely spin-degenerate bands[26, 25, 27, 28]. We identify two mechanisms for the generation of topological nodes in the electronic structure of CrSb. The first relies on the observation that spin-splitting of fully spin-polarized altermagnetic bands automatically leads to nodal lines/surfaces when opposite-spin bands cross. The degenerate states that form the nodal lines/surfaces have opposite spins. Such crossings in momentum space can be symmetry enforced, for instance at a mirror plane, or be accidental. The effect of SOC is to lift this spin-degeneracy for most, but not all momenta on the nodal line. At remaining WPs also the spin degeneracy remains, so that the points are a doublets with vanishing total spin-projection as illustrated in Fig. 1d. The altermagnetic symmetries and resulting spin-degeneracies thus provide a very natural setting for such opposite-spin WPs governed by the SOC energy scale. Indeed from DFT we identify 10 groups of opposite-spin WPs within 1 eV from the Fermi energy, see supplementary material (SM) for details.
@@
Interestingly, there is also a route that generates purely altermagnetic WPs, reflecting the bulk altermagnetic symmetries. It gives rise to same-spin WPs with a total spin-projection of +1/1+1/-1+ 1 / - 1 and a spin-splitting (between opposing spin projections) governed by the exchange interaction energy scale[1, 2], as illustrated in Fig. 1e. These points thus carry apart from a topological charge also a spin and thus a net magnetic moment. It is precisely the altermagnetic crystal symmetry that relates same-spin WPs with different spin-projections. These spin carrying WPs, and the topological surface states (SSs) they imply, reflect both the alternating nature of spins and the underlying relation between altermagnetism and Weyl semimetals.
@@
For CrSb, where lattice inversion symmetry connects atoms with the same magnetization, the SM provides details of the topological same-spin nodal line without SOC and three groups of same-spin WPs. Fig. 1f-1g shows the distribution of same-spin WPs in momentum space. Fermi-arcs connect the surface projections of these WPs of opposite chirality. As a consequence also their topological SSs are fully spin-polarized in the altermagnetic limit, with additional SOC affecting only details in their dispersion and spectral weight. The electronic structures of both Sb (Fig. 1h) and Cr (Fig. 1i) terminated surfaces show that the SS are very extended in momentum space and carry substantial spectral weight. Detailed representations of purely surface related states and the full 3D bulk band structures are provided in the SM.
@@
To experimentally access the spin-splitting and topological electronic structure of CrSb we use ARPES, which is to a certain extent complicated by CrSb having a 3D electronic structure. In order to establish the correspondence between photon energy and out-of-plane momentum kz, we performed broad-range (40 to 120 eV) photon energy dependent ARPES measurements along the $\rm\overline{M}-\overline{\Gamma}-\overline{M}$ direction. Fig. 2i shows the photon energy dependent ARPES spectral intensity map (kx-kz Fermi surface) at the Fermi level along the $\rm\overline{M}-\overline{\Gamma}-\overline{M}$ direction. From the periodic structure along the kz direction, the correspondence between the high symmetry points of the BZ along the kz direction and the photon energy is determined as shown in Fig. 2i. To distinguish between bulk and SSs, we measured Fermi surfaces on a relatively uneven sample surface for different kz, corresponding to photon energies from 97 eV to 122 eV (Fig. 2k). Since the uneven surface suppresses the contribution of SSs[39, 42], these Fermi surfaces mainly derive from the bulk. All their features are captured by the calculated bulk Fermi surface with corresponding kz (Fig. 2c-2h). The Fermi surfaces away from the kz = 0 $\pi/c$ and 1 $\pi/c$ planes (Fig. 2d-2f) have the form of a hexagram consisting of two intersecting equilateral triangular structures. Due to the altermagnetic order in CrSb, the two Fermi surfaces corresponding to the two different triangles have opposite-spin polarizations (Fig. 2d-2f), which a very clearly resolved experimentally. Interestingly, the Sb and Cr terminated surfaces (Fig. 2l and 2m) show additional features apart from the bulk Fermi surface, which are marked by orange arrows. We will show that these are related to the topological SSs resulting from the WPs.
@@
First we establish that there is a very large momentum dependent spin-splitting of the bulk bands. Fig. 3a-3c show the photon energy dependent dispersions along the $\rm\overline{M}-\overline{\Gamma}-\overline{M}$ direction measured with photon energies from 97 eV to 122 eV with the interval of 5 eV for an uneven sample surface (Fig. 3a) and flat Sb (Fig. 3b) and Cr (Fig. 3c) terminated surfaces. Along side it, the corresponding kz dependent calculated spin-polarized band structures are shown (Fig. 3d-3i). These bands are spin degenerate on the kz = 0 $\pi/c$ and 1 $\pi/c$ planes. The band structures between these two kz planes (Fig. 3e-3h) exhibit spin splitting due to the altermagnetism in CrSb. The calculated bands and their spin splittings are in very good agreement with the ARPES measurements [Fig. 3a-3c, (ii-v)]. We determine the maximum energy band splitting to be up to 1 eV (see Fig. S9 in the supplementary material). To further verify that the split bands are spin polarized, we performed spin resolved ARPES measurements with a photon energy of 102 eV, as shown in Fig. 3j-3l. Fig. 3j and 3k are the spin-polarized energy distribution curves (EDCs) along the left and right cut lines in Fig. 3a(ii). The spin-polarized intensity difference between spin-up and spin-down of left and right bands is shown in Fig. 3l. Due to the existence of a large non-spin polarized background, originating from a Shirley background, the measured spin polarization observed in the EDCs is limited. In spite of this, a difference in energy-dependent spin polarization intensity between spin-up and spin-down of left and right bands (Fig. 3l) can still be observed. This indicates that the left and right energy bands [Fig. 3a(ii)] have different spin projections along the $z$ direction, in agreement with theory.
@@
Having identified and quantified the altermagnetic spin-splitting in CrSb, we focus on the resulting topological properties. Both ARPES measurements and DFT calculations show that CrSb is metallic with a complex Fermi surface. In an energy range around from $-1$ eV to 1 eV around the Fermi level, we identify 13 groups of WPs in the calculated bands, as listed in Table. 1 of the SM. All WPs in CrSb arise from time-reversal symmetry breaking and their relative positions are given by the symmetries in the magnetic space group $P6_{3}^{\prime}/m^{\prime}m^{\prime}c$ (No. 194.268) reflecting their unique altermagnetic nature. Many WPs are along or near the $\rm\overline{\Gamma}-\overline{M}$ direction (see Fig. S2 in supplementary material) – this is the high-symmetry line whose little co-group allows non-zero chirality. The pair of WPs in the $\rm\overline{\Gamma}-\overline{M}$ direction below the Fermi level (marked by round dots in Fig. 4a-4b and 4h-4i) are same-spin WPs. Long SFAs connect neighboring WPs of opposite chirality and identical spin-projection across the BZ. This SS happens to appear in a region without bulk states so that it may be clearly discerned in ARPES. Even if the presence of these SFAs is dictated by topology, the shape of the surface spectra for two different terminations differ due to the lack of inversion and mirror symmetry on the surface, as shown in Fig. 4. Due to the large distance between two WPs projected on the 001 surface, the arc is also detectable along $\rm\overline{\Gamma}-\overline{K}$ as shown in Fig. 4o-4p.
@@
To track these WPs and SFAs experimentally, we conducted photon energy dependent band dispersion measurements along $\rm\overline{M}_{1}-\overline{\Gamma}-\overline{M}_{2}$ on Sb (Fig. 4c-4g) and Cr (Fig. 4j-4n) terminated surfaces. The band features measured on Sb (marked by orange arrows in Fig. 4c-4g) and Cr (marked by orange arrows in Fig. 4j-4n) surfaces exhibit negligible photon energy dependence [momentum distribution curve (MDC) analysis of band features measured on Cr terminated surface see Fig. S10 in SM], evidencing that it corresponds to a SS. Comparison with the corresponding calculated SSs, shows that their band features (marked by orange arrows in Fig. 4c-4g and Fig. 4j-4n) agree well with the SFAs (marked by orange arrows in Fig. 4b and 4i), indicating that the SFAs are intrinsic and robust. Their traces are also observed at the Fermi energy (Fig. 2l and 2m). Furthermore, the DFT calculations show the kz location of WPs at an energy of -0.357 eV (round dots in Fig. 4a-4b and 4h-4i) is $\pm$0.27 $\pi/c$. This corresponds to a photon energy of about 82 eV. We show the WP in the band dispersion panel measured with 82 eV (Fig. 4d) and indeed it is located right at the intersection of the bands. In addition, we also measured the band dispersion along $\rm\overline{\Gamma}-\overline{K}$ with photon energy of 85 eV on Sb (Fig. 4q) and Cr (Fig. 4t) surfaces (for more detailed photon energy dependent measurements along this direction see Fig. S11 in SM). Also along these high-symmetry lines, SFAs (denoted by orange arrows in Fig. 4q and 4t) can be observed for both surface terminations, reflecting the large portion of the surface BZ traversed by the SFA, consistent with the calculations.
Thus our high-resolution and spin-resolved ARPES measurements quantify the altermagnetic spin-spliting of bands in CrSb, establishing that it ranges from 1 eV to zero at high symmetry lines in the BZ. This momentum-dependent splitting leads to the presence of two types of topological nodal structures in the electronic structure, opposite-spin and same-spin WPs. The SFAs and WPs, as well as the consistency with their calculated properties, unequivocally establishes altermagnetic CrSb as a Weyl semimetal. Whereas opposite-spin WPs can also be encountered in canonical antiferromagnets, the spin-full points are beyond it. They produce spin-polarized SFAs with a magnetization that depends on the spin-projection of the WPs that they connect. The same-spin WPs carry two degrees of freedom: spin and topological charge. In altermagnets, crystal symmetries connecting the magnetic sublattices and improper rotation relate the same-spin WPs with different spin-projection and topological charge, respectively. These two degrees of freedom are locked, and consequently the physical response to external perturbations (such as strain and electric field) of spin will also reveal in topological features, particularly the ensuing topological SSs, and vice versa. This implies that the unique bulk altermagnetic spin-transport properties are promoted by topology to be properties of the Fermi arcs at the surface as well, which may render same-spin Weyl altermagnets interesting materials for spintronics interfaces. These insights not only spotlight the distinctive altermagnetic attributes of CrSb but also its capacity to induce novel physics and potential applications in the area of topological materials.

### v2

@@ Abstract
Altermagnets constitute a novel, third fundamental class of collinear magnetic ordered materials, alongside with ferro- and antiferromagnets [1, 2]. They share with conventional antiferromagnets the feature of a vanishing net magnetization. At the same time they show a spin-splitting of electronic bands, just as in ferromagnets, caused by the atomic exchange interaction[3, 4, 5]. On the other hand, topology has recently revolutionized our understanding of condensed matter physics, introducing new phases of matter classified by intrinsic topological order[6, 7, 8]. Here we connect the worlds of altermagnetism and topology, showing that the electronic structure of the altermagnet CrSb is topological and hosts a novel Weyl semimetallic state. Using high-resolution and spin angle-resolved photoemission spectroscopy, we observe a large momentum-dependent spin-splitting in CrSb, reaching up to 1 eV, that induces altermagnetic Weyl nodes with an associated magnetic quantum number. At the surface we observe their spin-polarized topological Fermi-arcs. This establishes that in altermagnets the large energy scale intrinsic to the spin-splitting – orders of magnitude larger than the relativistic spin-orbit coupling – creates its own realm of robust electronic topology.
@@
The recently established new class of altermagnetic materials is characterized by the presence of particular alternating opposite-spin sublattices. They defy conventional magnetic classifications by breaking time-reversal symmetry while simultaneously maintaining a zero net magnetization enforced by spin-lattice symmetry[1, 2]. Precisely this intrinsic trait allows altermagnets to combine advantageous features of both ferromagnets and antiferromagnets[9, 10], including spin-transport properties and compatibility with a diverse range of materials[11, 12, 13, 14, 15, 16, 17, 18], from superconductors[19, 20, 22, 23, 21] to insulators[24, 25, 26]. As the vanishing net magnetization in altermagnets avoids effects of long-range magnetic stray fields typical for ferromagnets, this combination opens up promising avenues for applications in magneto-optics[27], spintronics[28, 14], and beyond.
@@
Interestingly, symmetry analysis reveals that the electronic structure of altermagnets may in principle provide fertile ground for non-trivial electronic topology, due to the unique interplay between time-reversal and crystalline symmetries[30, 29, 31, 32, 33]. Despite extensive experimental and theoretical scrutiny of altermagnets[25, 35, 24, 26, 36, 37, 38, 34, 39, 40, 41] and topological semimetals[42, 6, 43, 44, 45, 46, 47, 48, 6, 43] separately, the simultaneous experimental manifestation of their characteristic phenomena has remained elusive so far. Here we establish CrSb as a topological Weyl semimetal with two distinct types of spin-carrying nodal structures. The giant spin splitting that we find from high resolution angle-resolved photoemission spectroscopy (ARPES) and spin-ARPES measurements aligns well with density functional theory (DFT) calculations on CrSb, as do the bulk Weyl points (WPs) with linear dispersion and ensuing altermagnetic surface Fermi arcs (SFAs) that, as we will show, connect same-spin WPs at the surface.
@@
For both its altermagnetic and topological features, the symmetries of CrSb are crucial. Figure 1a shows the crystal and magnetic structure of CrSb, which forms a hexagonal structure with the space group $P6_{3}/mmc$ (No. 194)[49]. The magnetic space group is $P6^{\prime}_{3}/m^{\prime}m^{\prime}c$ (No. 194.268) and the spin space group is $P^{-1}{6_{3}/}^{-1}{m}^{1}{m}^{-1}c^{\infty m}1$ with a six-fold screw rotation symmetry connecting Cr-atoms of opposite magnetic moment in real space. The three-dimensional (3D) Fermi surface of CrSb from DFT calculations in Fig. 1c illustrates that sixfold rotation connecting opposite-spin sublattices in real space (Fig. 1a) also connects opposite-spin electronic states in momentum space (Fig. 1c), as is mandatory for altermagnets[1, 2]. The calculated Fermi surface also illustrates the four mirror planes in the Brillouin zone (BZ) on which the altermagnetic bands are spin-degenerate in absence of spin-orbit coupling (SOC), as shown in Fig. 1b.
@@
The altermagnetic spin-splitting allows for topological electronic structures far beyond antiferromagnets with completely spin-degenerate bands[30, 29, 31, 32, 33]. We identify two mechanisms for the generation of topological nodes in the electronic structure of CrSb. The first relies on the observation that spin-splitting of fully spin-polarized altermagnetic bands automatically leads to nodal lines when opposite-spin bands cross. The degenerate states that form the nodal lines have opposite spins. Such crossings in momentum space can be symmetry enforced, for instance at a mirror plane, or be accidental. The effect of SOC in CrSb is to lift this spin-degeneracy for most, but not all momenta on nodal lines. At the remaining Weyl nodes also the spin degeneracy remains, so that the points correspond to doublets with vanishing total spin-projection $S^{z}=0$ as illustrated in Fig. 1d. The altermagnetic symmetries and resulting spin-degeneracies thus provide a very natural setting for such opposite-spin WPs governed by the SOC energy scale. Indeed from DFT we identify 10 groups of opposite-spin WPs within 1 eV from the Fermi energy, see supplementary material for details.
@@
Interestingly, there is also a route that generates purely altermagnetic WPs, reflecting the bulk altermagnetic symmetries. They give rise to same-spin WPs with a total spin-projection of $S^{z}=\pm 1$ (in units of $\hbar$) and a spin-splitting (energy-splitting at the nodal momentum) governed by the exchange interaction energy scale[1, 2], as illustrated in Fig. 1e. These altermagnetic Weyl nodes thus carry apart from a topological charge also a magnetic quantum number $S^{z}$, and with it, a finite magnetic moment. It is precisely the altermagnetic crystal symmetry that relates same-spin WPs with different $S^{z}$. The weak SOC that formally breaks the altermagnetic symmetry in CrSb causes the expectation value $|\langle S^{z}\rangle|$ to slightly deviate from unity (see supplemental material). These WPs, and the topological surface states they imply, render the topological features of Weyl altermagnets fundamentally distinct from ferro- or antiferromagnetic Weyl semimetals.
@@
For CrSb, where lattice inversion symmetry connects atoms with the same magnetization, the supplementary material provides details of the topological same-spin nodal line without SOC and three groups of same-spin WPs that we identified. Fig. 1f- 1g shows the distribution of same-spin WPs in momentum space, with their spin-projection quantum number. Fermi-arcs connect the surface projections of these WPs of opposite chirality. As a consequence also their topological surface states are fully spin-polarized in the altermagnetic limit, with additional SOC affecting the details of their dispersion and spectral weight. The electronic structures of both Sb (Fig. 1h) and Cr (Fig. 1i) terminated surfaces show that the surface states are very extended in momentum space and carry substantial spectral weight. Detailed representations of purely surface related states and the full 3D bulk band structures are provided in the supplementary material.
@@
To experimentally access the spin-splitting and topological electronic structure of CrSb we use ARPES, which is to a certain extent complicated by CrSb having a 3D electronic structure. In order to establish the correspondence between photon energy and out-of-plane momentum kz, we performed broad-range (40 to 120 eV) photon energy dependent ARPES measurements along the $\rm\overline{M}-\overline{\Gamma}-\overline{M}$ direction. Fig. 2i shows the photon energy dependent ARPES spectral intensity map (kx-kz Fermi surface) at the Fermi level along the $\rm\overline{M}-\overline{\Gamma}-\overline{M}$ direction. From the periodic structure along the kz direction, the correspondence between the high symmetry points of the BZ along the kz direction and the photon energy is determined as shown in Fig. 2i. To distinguish between bulk and surface states, we measured Fermi surfaces on a relatively uneven sample surface for different kz, corresponding to photon energies from 97 eV to 122 eV (Fig. 2k). Since the uneven surface suppresses the contribution of surface states[47, 50], these Fermi surfaces mainly derive from the bulk. All their features are captured by the calculated bulk Fermi surface with corresponding kz (Fig. 2c- 2h). The Fermi surfaces away from the kz = 0 $\pi/c$ and 1 $\pi/c$ planes (Fig. 2d- 2f) have the form of a hexagram consisting of two intersecting equilateral triangular structures. Due to the altermagnetic order in CrSb, the two Fermi surfaces corresponding to the two different triangles, which are very clearly resolved experimentally, have opposite-spin polarizations (Fig. 2d- 2f). Interestingly, the Sb and Cr terminated surfaces (Fig. 2l and  2m) show additional features apart from the bulk Fermi surface, which are marked by orange arrows. As we will detail later, comparison to the calculated surface spectral density (Fig. 2a- 2b) establishes that their root cause are the topological SFAs that in turn originate from the WPs (see Fig. S9 in supplementary material for a further comparison of constant energy contours).
@@
First we establish that there is a very large momentum dependent spin-splitting of the bulk bands in CrSb[1, 5, 36, 39, 40, 41]. Fig. 3a- 3c show the photon energy dependent dispersions along the $\rm\overline{M}-\overline{\Gamma}-\overline{M}$ direction measured with photon energies from 97 eV to 122 eV with the interval of 5 eV for an uneven sample surface (Fig. 3a) and flat Sb (Fig. 3b) and Cr (Fig. 3c) terminated surfaces. The corresponding kz dependent calculated spin-polarized band structures are shown alongside it (Fig. 3d- 3i). These bands are spin degenerate on the kz = 0 $\pi/c$ and 1 $\pi/c$ planes. Due to the altermagnetism in CrSb the bands between these two kz planes (Fig. 3e- 3h) exhibit spin splitting. The calculated bands and their spin splittings are in very good agreement with the ARPES measurements [Fig. 3a- 3c, (ii-v)]. We determine the maximum energy band splitting to be up to 1 eV (see Fig. S10 in the supplementary material). To further verify that the split bands are spin polarized, we performed spin resolved ARPES measurements with a photon energy of 102 eV, as shown in Fig. 3j- 3l. Fig. 3j and  3k are the spin-polarized energy distribution curves (EDCs) along the left and right cut lines in Fig. 3a(ii). The spin-polarized intensity difference between spin-up and spin-down of left and right bands is shown in Fig. 3l. Due to the existence of a large non-spin polarized background, originating from a Shirley background, the measured spin polarization observed in the EDCs is limited. Despite this, a difference in energy-dependent spin polarization intensity between spin-up and spin-down of left and right bands (Fig. 3l) can still be observed. This indicates that the left and right energy bands [Fig. 3a(ii)] have different spin projections along the $z$ direction, in agreement with theory.
@@
Having identified and quantified the altermagnetic spin-splitting in CrSb, we focus on the resulting topological properties. Both ARPES measurements and DFT calculations show that CrSb is metallic with a complex Fermi surface. In an energy range around from $-1$ eV to 1 eV around the Fermi level, we identify 13 groups of WPs in the calculated bands, as listed in Table. 1 of the supplementary material. All WPs in CrSb arise from time-reversal symmetry breaking and their relative positions are given by the symmetries in the magnetic space group $P6_{3}^{\prime}/m^{\prime}m^{\prime}c$ (No. 194.268) reflecting their unique altermagnetic nature. Many WPs are along or near the $\rm\overline{\Gamma}-\overline{M}$ direction (see Fig. S2 in supplementary material) – this is the high-symmetry line whose little co-group allows non-zero chirality. The pair of WPs in the $\rm\overline{\Gamma}-\overline{M}$ direction below the Fermi level (marked by round dots in Fig. 4a- 4b and  4h- 4i) are same-spin WPs. Long SFAs connect neighboring WPs of opposite chirality $\chi=\pm 1$ and identical spin-projection across the BZ. This topological surface state happens to appear in a region without bulk states so that it may be clearly discerned in ARPES. Even if the presence of these SFAs is dictated by topology, the shape of the surface spectra for two different terminations differ due to the lack of inversion and mirror symmetry on the surface, as shown in Fig. 4. Due to the large distance between two WPs projected on the 001 surface, the arc is even detectable along $\rm\overline{\Gamma}-\overline{K}$ as shown in Fig. 4o- 4p.
@@
To track these WPs and SFAs experimentally, we conducted photon energy dependent ARPES measurements along $\rm\overline{M}_{1}-\overline{\Gamma}-\overline{M}_{2}$ on Sb (Fig. 4c- 4g) and Cr (Fig. 4j- 4n) terminated surfaces. The band features measured on Sb (marked by orange arrows in Fig. 4c- 4g) and Cr (marked by orange arrows in Fig. 4j- 4n) surfaces exhibit negligible photon energy dependence [momentum distribution curve (MDC) analysis of band features measured on Cr terminated surface see Fig. S11 in supplementary material], evidencing that it corresponds to a surface state. Comparison with the corresponding calculated surface states, shows that their band features (marked by orange arrows in Fig. 4c- 4g and Fig. 4j- 4n) agree well with the SFAs (marked by orange arrows in Fig. 4b and  4i), indicating that the SFAs are intrinsic and robust. As pointed out earlier, the SFAs are also observed in the 2D momentum maps at the Fermi energy (Fig. 2l and  2m). Furthermore, the DFT calculations show the kz location of WPs at an energy of -0.357 eV (round dots in Fig. 4a- 4b and  4h- 4i) is $\pm$0.27 $\pi/c$. This corresponds to a photon energy of about 82 eV. We show the WP in the band dispersion panel measured with 82 eV (Fig. 4d) and indeed it is located right at the intersection of the bands. In addition, we also measured the band dispersion along $\rm\overline{\Gamma}-\overline{K}$ with photon energy of 85 eV on Sb (Fig. 4q) and Cr (Fig. 4t) surfaces (for more detailed photon energy dependent measurements along this direction see Fig. S12 in supplementary material). Also along these high-symmetry lines, SFAs (denoted by orange arrows in Fig. 4q and  4t) are observed for both surface terminations, reflecting the large portion of the surface BZ traversed by the SFA, consistent with the calculations.
Thus we have not only quantified the altermagnetic spin-spliting of bands in CrSb with high-resolution and spin-resolved ARPES measurements, but also established it as a topological Weyl semi-metal, with concomitant altermagnetic topological surface states. Whereas in general Weyl nodes carry only a topological charge, corresponding to their chirality $\chi$, in altermagnets they also carry spin-projection $S^{z}$. Besides opposite-spin nodes with zero spin-projection, altermagnetic CrSb also has same-spin nodes with spin-projection $\pm\hbar$ and chirality-spin locking. For each set of $n$ symmetry-related altermagnetic Weyl nodes one can define a locking invariant $\zeta=\sum_{i=1,n}\chi_{i}S^{z}_{i}$. For CrSb $\zeta=0$ by symmetry, but interestingly other magnetic symmetry groups allow sets of altermagnetic nodes with non-zero integer locking invariants. The distribution of same-spin nodes in 3D momentum space reflects all altermagnetic symmetries as do their spin-polarized topological surface states, the altermagnetic Fermi arcs. Spin-orbit coupling weakly breaks altermagnetic symmetries and consequently the quantization of the spin-projection. As these properties derive from the interplay of symmetry and topology, they are generic for Weyl altermagnets. These findings may well imply that the unique bulk altermagnetic spin-transport properties are promoted by topology to also become properties of the Fermi arcs at the surface, which can render same-spin Weyl altermagnets interesting interface materials for spintronics. These insights not only spotlight the distinctive altermagnetic attributes of CrSb but also its potential to induce novel physics and applications in the area of topological materials.